Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  RH-ISAC, Page 45 of 141  Not logged in ELOG logo
ID Date Author Category Type Module Target/Number Subject Text Attachments
  1926   Wednesday, May 22, 2019, 08:33 Travis CaveConditioning StationStandard OperationTM1no targetModule move TM#1 has been moved from the north
west silo to the conditioning station for
the day it will be moved back to the silo
  
  1925   Friday, May 17, 2019, 16:46 Adam NewsomeSafe Module ParkingDevelopment  SMP Controls Installation Update 
The following controls-related
 IMG_20190417_100838.jpgIMG_20190417_100843.jpgIMG_20190517_162230.jpgIMG_20190517_162305.jpg 
  1924   Friday, May 17, 2019, 10:47 Travis CaveITEStandard OperationTM2Ta#58Module move TM#2 with Ta#58 has been moved from
the south hot cell to the east target station.
  
  1923   Thursday, May 16, 2019, 14:35 Travis CaveSouth Hot-CellStandard OperationTM2Ta#58Electrical Check Results of the electrical check of
TM#2 without target and with the new Ta#58
target.
 Ta#58_TM2_SHC_electrical_chk_May_16_2019.pdfElectrical_Chk_TM2_no_target_May_15_2019_SHC.pdf 
  1922   Thursday, May 16, 2019, 12:31 chad fisherSouth Hot-CellStandard OperationTM2Ta#58Ta #58 installedTa #58 has been installed onto TM2. All
new silver plated high current fasteners.
All torqued to 130 in lbs.
  
  1921   Wednesday, May 15, 2019, 16:06 Isaac EarleSafe Module ParkingDevelopment  SMP module support flange installed and leak checkedThe SMP module support flange sub-assembly
(IRH1755) was installed today according to
drawing IRH1683.  Dow Corning 111 sealing
 20190515_094730.jpg20190515_094740.jpg20190515_103253.jpg20190515_103824.jpg 
  1920   Monday, May 13, 2019, 11:22 David WangITWRepair  ITW entrance module has been connnected.Entrance module has been connected back
to ITW. Everything is normal.
  
  1919   Monday, May 13, 2019, 09:55 Travis CaveITWStandard Operation  Entrance module move The West Entrance module has been
moved from the south hot cell to the west
target station.
  
  1918   Friday, May 10, 2019, 10:23 chad fisherSouth Hot-CellRepair  West Entrance module wire scanner replacementWest entrance module - visible broken or
disconnected wire visible on old/existing
wire scanner. Wire scanner removed, new one
 20190509_085908.jpg20190510_101235.jpg 
  1917   Wednesday, May 08, 2019, 10:29 Travis CaveSouth Hot-CellStandard Operation  Entrance module move The entrance module has been moved
from the west station to the south hot cell.
  
  1916   Thursday, May 02, 2019, 09:37 Travis CaveSpent Target VaultStandard Operation UC#26Spent Target Move The spent UC#26 target was moved
from the south hot cell to the spent target
vault. It was place in pail #209 and put
 Target_Index_2019-05-02.pdf 
  1915   Wednesday, May 01, 2019, 10:45 Travis CaveSouth Hot-CellStandard OperationTM4no targetElectrical Check Electrical check of the TM#4 with
no target attached at the south hot cell
before being moved to the south west silo.
 TM4_elec_chk_no_tgt_shc_may_1_2019.pdf 
  1914   Wednesday, May 01, 2019, 09:34 Travis CaveSouth Hot-CellStandard OperationTM4no targetModule move TM#4 has now been moved to the south
west silo.
  
  1913   Wednesday, May 01, 2019, 07:49 chad fisherSouth Hot-CellStandard OperationTM4UC#26UC#26 removalUC#26 has been removed from TM4.  20190501_072940[1].jpg 
  1912   Tuesday, April 30, 2019, 11:19 Travis CaveSouth Hot-CellStandard OperationTM4UC#26Electrical Check Electrical check of TM#4 with UC#26
post beam at the south hot cell. See attached.
 Elec_CHk_UC26_SHC_post_beam.pdf 
  1911   Tuesday, April 30, 2019, 08:59 Travis CaveSouth Hot-CellStandard OperationTM4UC#26Module move TM#4 with UC#26 has been moved from
ITE to the south hot cell.
  
  1910   Monday, April 29, 2019, 09:32 Travis CaveSpent Target VaultStandard Operation SiC#31,32 and UC#19,20,21Spent targets moved to mini vault. The following targets have been moved
from the spent target vault to the mini vault:
#127 - SiC#31 (front)
  
  1909   Monday, April 29, 2019, 07:56 chad fisherSouth Hot-CellDevelopment  Pail transfersTwo pail transfers took place April 25th.
Ta#47 transfered from pail 137 to
pail 201. existing plastic bag from 137 showed
 20190425_133750.jpg20190425_134106.jpg 
  1908   Thursday, April 18, 2019, 10:04 Travis CaveSouth Hot-CellMaintenance Ta#50, SiC#34, Ta#46, Ta#47, and Ta#45Target transfer from old pail to new pail. Targets Ta#50 and SiC#34 have been
transferred from their old pails to new ones.
Ta#50 was done on April 12th 2019 it was
  
  1907   Wednesday, April 17, 2019, 12:03 chad fisherSouth Hot-CellRepair  Manipulator repairBroken handle cable of east manipulator
has been replaced and manipulator is operational
again.
  
ELOG V2.9.2-2455